OpenStudy (bossimbacon):

Calib received $50 for his birthday and put it in his piggy bank. Each week, he puts 2 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 2w + 50 where w is the number of weeks he has saved 2 dollars. His sister Lara makes a table of reasonable values for w. Lara selects 14 as a value for w. What is the corresponding table entry for D? A. 66 B. 78 C. 57 D. 264 @Destinymasha

OpenStudy (anonymous):

Plug 14 in W D=2W+50 D=2(14)+50 D=28+50 D=78